# The $S_{n}$-equivariant Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$

### v1

@@ 5. Calculation of $\mathsf{Dih}_{r}$
$\mathsf{e}(\mathsf{Dih}_{r})$ $=\sum_{d\geq 2}q^{d}\left(\sum_{k=2}^{d}\eta_{k,d}(r)\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\sum_{j\mid k}\theta_{j,k,d}(r)\mathfrak{p}_{j}^{k/j}\right).$
[…]
@@
The relevance of these definitions is as follows. Let $O_{k}$ denote a set of orbit representatives for the action of $D_{k}$ on its subgroups by conjugation. Grouping $[\delta]\in\Gamma_{r,d}(k)$ according to the conjugacy class types of $D_{k}(\delta)$, we have
@@
[…]
@@ 5.1. Conjugacy classes of dihedral subgroups
The subgroup structure of $D_{k}$ is well-known, see e.g. [Con19, §3].
@@
If $k$ and $j$ are even while $c$ is odd, then $[\langle\rho^{j},\rho^{c}\tau\rangle]=[\langle\rho^{j^{\prime}},\rho\tau\rangle]$ where $j^{\prime}$ is even. Thus we focus on computing $\gamma^{[\langle\rho^{j}\rangle]}_{r,d}(k)$ for all $k$ and $j\mid k$, as well as the numbers $\gamma^{[\langle\rho^{j},\rho\tau\rangle]}_{r,d}(k)$ when $k$ is even and $j\mid k$ is an even divisor.
@@ 5.3. Möbius inversion
The second two cases follow from the fact that the intervals $[\langle\rho^{a}\rangle,\langle\rho^{b}\rangle]$ and $[\langle\rho^{a},\rho^{c}\tau\rangle,\langle\rho^{b},\rho^{c}\tau\rangle]$ are each isomorphic as lattices to the lattice of divisors of $a/b$. For the first case, we argue as follows. If $G$ is an arbitrary finite group, let $\mu_{G}$ denote the Möbius function of the poset $\mathcal{L}(G)$ of subgroups of $G$. If $N$ is a normal subgroup of $G$ which is contained in another subgroup $H$ of $G$. Then
@@
[…]
@@
We have computed $\mu_{D_{s}}(1,\langle\rho^{i}\rangle)=\mu(s/i)$. For each divisor $i|s,i\neq 1$, we have $\langle\rho^{i},\rho^{j}\tau\rangle\cong D_{s/i}$, so from inductive hypothesis there is $\mu_{D_{s}}(1,\langle\rho^{i},\rho^{j}\tau\rangle)=-(s/i)\mu(s/i)$. Summing up, we have
@@ 5.4. Quotient graphs
We enumerate the sets $\tilde{\Gamma}_{r,d}(k)^{H^{\prime}}$ via the following construction of quotient graphs. This is an established technique for enumerating graphs with automorphisms [Han85, §1].
@@
[…]
@@
[…]
@@ 5.5. Strata enumeration
denote the chromatic polynomial of $C_{k}$ [Bir12, Whi32]. They yield straightforward formulas for the numbers $\tilde{\gamma}_{r,d}(k)=|\tilde{\Gamma}_{r,d}(k)|$ and $|\tilde{\varepsilon}_{r,d}(k)|$.
@@
$\left|\tilde{\Gamma}_{r,d}(k)^{\langle\rho^{2\ell},\rho^{c}\tau\rangle}\right|=\binom{\frac{d\ell}{k}-1}{\ell-1}\omega_{k}(r+1).$
@@ 5.6. Computing the characters
In both cases we have $\mathsf{e}^{2\wr k}(\operatorname{Ind}_{H}^{B_{k}}\operatorname{Spec}\mathbb{C})=\operatorname{ch}_{2\wr k}(\operatorname{Ind}_{H}^{B_{k}}\mathrm{Triv}_{H}).$ As explained by Petersen [Pet12],
@@
$\mathsf{e}(\mathsf{Dih}_{r})$ $=\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\sum_{j\mid k}\frac{G_{j,k,d}(r)}{4}\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\left(\frac{F_{j,k,d}(r)}{2k}-\frac{jG_{j,k,d}(r)}{4k}+\frac{jG_{j,k,d}(r)}{4k}\right)\sum_{s\mid\frac{k}{j}}\varphi(s)\mathfrak{p}_{s}^{k/s}$
$=\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\sum_{j\mid k}\frac{G_{j,k,d}(r)}{4}\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\frac{F_{j,k,d}(r)}{2k}\sum_{s\mid\frac{k}{j}}\varphi(s)\mathfrak{p}_{s}^{k/s}$
$=\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\eta_{k,d}(r)\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\sum_{s\mid k}\mathfrak{p}_{s}^{k/s}\sum_{j\mid\frac{k}{s}}\frac{\varphi(s)}{2k}F_{j,k,d}(r)$
$=\sum_{d\geq 0}q^{d}\sum_{k=2}^{d}\eta_{k,d}(r)\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\sum_{s\mid k}\theta_{s,k,d}(r)\mathfrak{p}_{s}^{k/s}$
@@ 6. Calculation of $\mathscr{B}_{r}$
The Serre characteristic $\mathsf{e}(\mathsf{Dih}_{r})$ was computed in §5. In this section, we will explain a formula for $\mathsf{e}(\mathsf{Cat})$ that appears in [BM13, Remark 4.2], and then compute the plethysm $\mathsf{e}(\mathsf{Dih}_{r})\circ_{S_{2}}\mathsf{e}(\mathsf{Cat})$. Combining this with Getzler’s calculation of $\mathscr{B}_{0}$ will give our main technical Theorem B, recalled below, which implies Theorem A.
@@
$\phantom{space}+\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\eta_{k,d}(r)\frac{(1+2\dot{\mathscr{A}}_{0})^{2}}{(1-\psi_{2}(\mathscr{A}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathscr{A}_{0}))^{k/j}}.$
@@
given by Bergström–Minabe [BM13]. The formula, recalled in Theorem 6.2 below, expresses $\mathsf{e}(\mathsf{Cat})$ in terms of the known series
@@
The formula is stated in [BM14, Remark 4.2]. Here we explain how it follows from their main result on Serre characteristics of the Losev–Manin spaces ¯$\overline{\mathcal{M}}_{0,2|n}$. In Proposition 4.1 of loc. lit., they prove that
@@
The modular interpretation of $\mathsf{Cat}$ implies that its characteristic is obtained from $\sum_{n}\mathsf{e}^{S_{2}\times S_{n}}(\overline{\mathcal{M}}_{0,2|n})$ via ‘distinguishing’ the light marked points, i.e., by applying the plethystic logarithm as in [GK98, Proposition 8.6]:
@@
where the first equality follows from [KSY24, Proposition 4.3], and the second follows from Lemma 2.12. From the way $f_{i}$ and $g_{i}$ are defined, it follows that
@@
$\mathsf{e}(\mathsf{Dih}_{r})\circ_{S_{2}}\mathsf{e}(\mathsf{Cat})=\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\eta_{k,d}(r)\frac{(1+2\dot{\mathscr{A}}_{0})^{2}}{(1-\psi_{2}(\mathscr{A}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathscr{A}_{0}^{\prime\prime}))^{k/j}}.$
@@
We now conclude by indicating how Theorems A and B follows from Theorem 6.4: in [Get98] (with a typo corrected in [Pet12, Page 1]), Getzler proves
@@
7. Calculations and applications
Calculations of $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d))$ for small $d$ and $n$, performed in SageMath using Theorem A, are found at the end of this section in Tables 1, 2, and 3. In each case, by treating $r$ as a formal variable, we see that the multiplicity of each irreducible representation of $S_{n}$ in $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d))$ is equal to an integral linear combination of binomial coefficients of the form $\binom{r+1}{k}$, for $1\leq k\leq d+1$. We now state Theorem C and its corollaries, which imply that this pattern holds in arbitrary genus, on the level of virtual $S_{n}$-equivariant mixed Hodge structures. The proof of Theorem C is after the statement of Lemma 7.3. Further discussion of the included data follows, in §7.1 and §7.2.
Theorem C.
There exists a polynomial $\mathsf{K}_{g,n,d}(t)\in(K_{0}(\mathsf{M})\otimes\Lambda)[t]$ of degree $d$ in $t$ such that
$\mathsf{e}^{S_{n}}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}})=(r+1)\cdot\mathsf{K}_{g,n,d}(r+1)$
for all integers $r>0$.
Recall that for an $S_{n}$-variety $X$, the invariant $\mathsf{e}_{\lambda}(X)\in K_{0}(\mathsf{M})$ defined in (2.7) records the multiplicity of the Specht module $V_{\lambda}$ in $\mathsf{e}^{S_{n}}(X)$, while $\chi_{\lambda}(X)\in\mathbb{Z}$ is the rank of $\mathsf{e}_{\lambda}(X)$. The following result is an immediate corollary of Theorem C.
Corollary 7.1.
For each $\lambda\vdash n$, there exists a unique polynomial $m_{g,n,d}^{\lambda}(t)\in(K_{0}(\mathsf{M})\otimes\mathbb{Q})[t]$ of degree at most $d$ such that
$\mathsf{e}_{\lambda}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}})=(r+1)\cdot m^{\lambda}_{g,n,d}(r+1)$
for all integers $r>0$.
By definition, the polynomial $m_{g,n,d}^{\lambda}(t)$ of Corollary 7.1 above must be “integer-valued” in the sense that evaluation at integers produces elements of $K_{0}(\mathsf{M})$ rather than $K_{0}(\mathsf{M})\otimes\mathbb{Q}$. This observation, together with Pólya’s theorem on integer-valued polynomials [Pól15], gives the following corollary.
Corollary 7.2.
For each $\lambda\vdash n$, there is an integer-valued function
$a_{g,d,\lambda}:\{1,\ldots,d+1\}\to\mathbb{Z}$
such that
$\chi_{\lambda}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d))=\sum_{k=1}^{d+1}a_{g,d,\lambda}(k)\binom{r+1}{k}$
for all integers $r>0$.
Theorem C is a consequence of basic properties of the unlabeled chromatic polynomial studied by Hanlon [Han85] and our $S_{n}$-equivariant formulation (4.5) of localisation for ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$. For an arbitrary graph $G$, let $\mathrm{Col}_{G}(x)$ denote the set of $x$-colourings of $G$. Given a subgroup $H\leq G$, set $\mathrm{Col}_{G}^{[H]}(x)\subset\mathrm{Col}_{G}(x)$ for the subset of $x$-colourings whose stabiliser in $\operatorname{Aut}(G)$ is conjugate to $H$. Then the $\operatorname{Aut}(G)$-action on $\mathrm{Col}_{G}(x)$ preserves $\mathrm{Col}_{G}^{[H]}(x)$, and we define
$P^{[H]}_{G}(x):=\left|\frac{\mathrm{Col}_{G}^{[H]}(x)}{\operatorname{Aut}(G)}\right|.$
The following lemma is a straightforward consequence of [Han85, Corollary 2.1], together with Möbius inversion.
Lemma 7.3.
[…]
With Lemma 7.3, we can prove Theorem C.
Proof of Theorem C.
For an integer $r>0$, consider the functor
$\alpha:\widetilde{\mathsf{LG}}_{g,n,d}(r)\to\widetilde{\mathsf{LG}}_{g,n,d}(0)$
given by forgetting the coloring. Given a localisation graph $\mathcal{G}$ in $\widetilde{\mathsf{LG}}_{g,n,d}(0)$, we write $\mathsf{C}_{\mathcal{G}}(r+1)$ for the fibre of $\alpha$ over $\mathcal{G}$, which is the groupoid of $(\mathbb{P}^{r})^{\mathbb{C}^{\star}}$-colorings of the decorated graph $\mathcal{G}$. Applying equations (4.4) and (4.5) as well as Lemma 3.1, we then have $S_{n}$-equivariant isomorphisms
¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ $\cong\coprod_{\mathcal{G}\in\pi_{0}\left(\widetilde{\mathsf{LG}}_{g,n,d}(0)\right)}\coprod_{c\in\pi_{0}(\mathsf{C}_{\mathcal{G}}(r+1))}\left(\operatorname{Ind}_{\prod_{v\in V(\mathcal{G})}S_{n(v)}}^{S_{n}}\overline{\mathcal{M}}_{g(v),n(v)+\mathrm{val}(v)}\right)/\operatorname{Aut}(\mathcal{G},c)$
$\cong\coprod_{\mathcal{G}\in\pi_{0}\left(\widetilde{\mathsf{LG}}_{g,n,d}(0)\right)}\coprod_{c\in\pi_{0}(\mathsf{C}_{\mathcal{G}}(r+1))}\left(\operatorname{Ind}_{\operatorname{Aut}(\mathcal{G},c)}^{\operatorname{Aut}(\mathcal{G})}\operatorname{Ind}_{\prod_{v\in V(\mathcal{G})}S_{n(v)}}^{S_{n}}\overline{\mathcal{M}}_{g(v),n(v)+\mathrm{val}(v)}\right)/\operatorname{Aut}(\mathcal{G}),$
[…]
$Z_{\mathcal{G},[H]}:=\left[\left(\operatorname{Ind}_{H}^{\operatorname{Aut}(\mathcal{G})}\operatorname{Ind}_{\prod_{v\in V(\mathcal{G})}S_{n(v)}}^{S_{n}}\overline{\mathcal{M}}_{g(v),n(v)+\mathrm{val}(v)}\right)/\operatorname{Aut}(\mathcal{G})\right]\in K_{0}(\mathsf{V},\mathbb{S}),$
where $H\in[H]$ acts as a subgroup of $\operatorname{Aut}(\mathcal{G})$. We observe that the number of isomorphism classes $c\in\pi_{0}(\mathsf{C}_{\mathcal{G}}(r+1))$ with $[\operatorname{Aut}(\mathcal{G},c)]=[H]$ is equal to $P_{\mathcal{G}}^{[H]}(r+1)$. Thus in $K_{0}(\mathsf{V},\mathbb{S})$ we have an equality
[¯$[\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}]=\sum_{\mathcal{G}\in\pi_{0}\left(\widetilde{\mathsf{LG}}_{g,n,d}(0)\right)}\sum_{[H]\leq G}P_{\mathcal{G}}^{[H]}(r+1)Z_{\mathcal{G},[H]},$
So the Theorem now follows from Lemma 7.3 and the observation that $|V(\mathcal{G})|\leq d+1$ for any $d$ localisation graph $\mathcal{G}$. ∎
7.1. Calculation of $\mathsf{b}_{1,r}^{\mathrm{nrt}}$
We now discuss the tables of data at the end of the paper. The first step in producing Tables 1, 2, and 3 for is the explicit calculation of $\mathsf{b}_{1,r}^{\mathrm{nrt}}$ in low degrees. The calculation is particularly accessible when $n=0$. In this case, the fixed locus
¯$\overline{\mathcal{M}}_{1,0}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$
[…]
Corollary 7.4.
For $d\geq 2$,
$\mathsf{e}\left(\overline{\mathcal{M}}_{1,0}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}\right)=\sum_{k=2}^{d}\eta_{k,d}(r)+\sum_{j\mid k}\theta_{j,k,d}(r).$
In general, to compute $\mathsf{b}_{1,r}^{\mathrm{nrt}}$, we recall the following formula from Theorem 4.1:

### v2

@@ 5. Calculation of $\mathsf{Dih}_{r}$
$\mathsf{e}(\mathsf{Dih}_{r})$ $=\sum_{d\geq 2}q^{d}\left(\sum_{k=2}^{d}\left(\eta_{k,d}(r)\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\sum_{j\mid k}\theta_{j,k,d}(r)\mathfrak{p}_{j}^{k/j}\right)\right).$
The proof of Theorem 5.5 amounts to enumerating localisation graphs with fixed isomorphism type of automorphism group. Given $\delta]\in\Gamma_{r,d}(k)$, a choice of lift $\delta$ determines a subgroup $D_{k}(\delta)$ of $D_{k}$, well-defined up to conjugacy. Given a conjugacy class of subgroups $[H]$ of $D_{k}$, we let
@@
The relevance of these definitions is as follows. Let $O_{k}$ denote a set of orbit representatives for the action of $D_{k}$ on its subgroups by conjugation. Grouping $\delta]\in\Gamma_{r,d}(k)$ according to the conjugacy class types of $D_{k}(\delta)$, we have
@@
where $H$ is any choice of representative of $[H]$. We will spend the rest of this section explaining how to understand the set $O_{k}$ and giving formulas for $\gamma^{[H]}_{r,d}(k)$ as $[H]$ varies over all elements of $O_{k}$.
@@ 5.1. Conjugacy classes of dihedral subgroups
The subgroup structure of $D_{k}$ is well-known, see e.g. [15, §3].
@@
If $k$ and $j$ are even while $c$ is odd, then $\langle\rho^{j},\rho^{c}\tau\rangle]=[\langle\rho^{j^{\prime}},\rho\tau\rangle]$ where $j^{\prime}$ is even. Thus we focus on computing $\gamma^{[\langle\rho^{j}\rangle]}_{r,d}(k)$ for all $k$ and $j\mid k$, as well as the numbers $\gamma^{[\langle\rho^{j},\rho\tau\rangle]}_{r,d}(k)$ when $k$ is even and $j\mid k$ is an even divisor.
@@ 5.3. Möbius inversion
The second two cases follow from the fact that the intervals
$\langle\rho^{a}\rangle,\langle\rho^{b}\rangle]\quad\mbox{and}\quad[\langle\rho^{a},\rho^{c}\tau\rangle,\langle\rho^{b},\rho^{c}\tau\rangle]$
are each isomorphic as lattices to the lattice of divisors of $a/b$. For the first case, we argue as follows. If $G$ is an arbitrary finite group, let $\mu_{G}$ denote the Möbius function of the poset $\mathcal{L}(G)$ of subgroups of $G$. If $N$ is a normal subgroup of $G$ which is contained in another subgroup $H$ of $G$. Then
@@
since the intervals $[H,G]$ in $\mathcal{L}(G)$ and $[H/N,G/N]$ in $\mathcal{L}(G/N,H/N)$ are isomorphic as lattices. Since cyclic subgroups of $D_{k}$ are normal, we have
@@
We have computed $\mu_{D_{s}}(1,\langle\rho^{i}\rangle)=\mu(s/i)$. For each divisor $i|s,i\neq 1$, we have $\langle\rho^{i},\rho^{j}\tau\rangle\cong D_{s/i}$, so from the inductive hypothesis there is $\mu_{D_{s}}(1,\langle\rho^{i},\rho^{j}\tau\rangle)=-(s/i)\mu(s/i)$. Summing up, we have
@@ 5.4. Quotient graphs
We enumerate the sets $\tilde{\Gamma}_{r,d}(k)^{H^{\prime}}$ via the following construction of quotient graphs. This is an established technique for enumerating graphs with automorphisms [33, §1].
@@
We write $\tilde{\varepsilon}_{r,d}(k)$ for the set of $(r,d)$-decorations of a labelled path $P_{k}$ with $k+1$ ordered vertices; formally, such a decoration is an $(r+1)$-colouring $\delta_{V}:V(P_{k})\to\{0,\ldots,r+1\}$ of $P_{k}$, together with a function $\delta_{E}:E(P_{k})\to\mathbb{Z}_{>0}$ with $\sum_{e\in E(P_{k})}\delta_{E}(e)=d$.
@@
Inverses of the maps are given by pulling back vertex colourings and degree decorations on $C_{k,H}$ along the quotient map of vertices and edges. One can check that the pullback does define a $(r,d)$-decoration on $C_{k}$. Therefore, the pairs of sets are in bijection. ∎
@@ 5.5. Strata enumeration
denote the chromatic polynomial of $C_{k}$ [11, 67]. They yield straightforward formulas for the numbers $\tilde{\gamma}_{r,d}(k)=|\tilde{\Gamma}_{r,d}(k)|$ and $|\tilde{\varepsilon}_{r,d}(k)|$.
@@
$\left|\tilde{\Gamma}_{r,d}(k)^{\langle\rho^{2\ell},\rho^{c}\tau\rangle}\right|=\binom{\frac{d\ell}{k}-1}{\ell-1}\omega_{\ell}(r+1).$
@@ 5.6. Computing the characters
In both cases we have $\mathsf{e}^{2\wr k}(\operatorname{Ind}_{H}^{B_{k}}\operatorname{Spec}\mathbb{C})=\operatorname{ch}_{2\wr k}(\operatorname{Ind}_{H}^{B_{k}}\mathrm{Triv}_{H}).$ As explained in [54],
@@
$\mathsf{e}(\mathsf{Dih}_{r})$ $=\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\sum_{j\mid k}\left(\frac{G_{j,k,d}(r)}{4}\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\left(\frac{F_{j,k,d}(r)}{2k}-\frac{jG_{j,k,d}(r)}{4k}+\frac{jG_{j,k,d}(r)}{4k}\right)\sum_{s\mid\frac{k}{j}}\varphi(s)\mathfrak{p}_{s}^{k/s}\right)$
$=\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\sum_{j\mid k}\left(\frac{G_{j,k,d}(r)}{4}\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\frac{F_{j,k,d}(r)}{2k}\sum_{s\mid\frac{k}{j}}\varphi(s)\mathfrak{p}_{s}^{k/s}\right)$
$=\sum_{d\geq 2}q^{d}\left(\sum_{k=2}^{d}\eta_{k,d}(r)\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\sum_{s\mid k}\left(\mathfrak{p}_{s}^{k/s}\sum_{j\mid\frac{k}{s}}\frac{\varphi(s)}{2k}F_{j,k,d}(r)\right)\right)$
$=\sum_{d\geq 0}q^{d}\sum_{k=2}^{d}\left(\eta_{k,d}(r)\mathfrak{p}_{2}^{k/2-1}\mathfrak{q}_{1}^{2}+\left(\sum_{s\mid k}\theta_{s,k,d}(r)\mathfrak{p}_{s}^{k/s}\right)\right)$
@@ 6. Calculation of $\mathscr{B}_{r}$
The Serre characteristic $\mathsf{e}(\mathsf{Dih}_{r})$ was computed in §5. In this section, we will explain a formula for $\mathsf{e}(\mathsf{Cat})$ that appears in [7, Remark 4.2], and then compute the plethysm $\mathsf{e}(\mathsf{Dih}_{r})\circ_{S_{2}}\mathsf{e}(\mathsf{Cat})$. Combining this with Getzler’s calculation of $\mathscr{B}_{0}$ will give our main technical Theorem B, recalled below, which implies Theorem A.
@@
$\phantom{space}+\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\left(\eta_{k,d}(r)\frac{(1+2\dot{\mathscr{A}}_{0})^{2}}{(1-\psi_{2}(\mathscr{A}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathscr{A}_{0}^{\prime\prime}))^{k/j}}\right).$
@@
given by Bergström–Minabe [7]. The formula, recalled in Theorem 6.2 below, expresses $\mathsf{e}(\mathsf{Cat})$ in terms of the known series
@@
The formula is stated in [8, Remark 4.2]. Here we explain how it follows from their main result on Serre characteristics of the Losev–Manin spaces $\overline{\mathcal{M}}_{0,2|n}$. In Proposition 4.1 of loc. lit., they prove that
@@
The modular interpretation of $\mathsf{Cat}$ implies that its characteristic is obtained from $\sum_{n}\mathsf{e}^{S_{2}\times S_{n}}(\overline{\mathcal{M}}_{0,2|n})$ via ‘distinguishing’ the light marked points, i.e., by applying the plethystic logarithm as in [21, Proposition 8.6]:
@@
where the first equality follows from [35, Proposition 4.3], and the second follows from Lemma 2.12. From the way $f_{i}$ and $g_{i}$ are defined, it follows that
@@
$\mathsf{e}(\mathsf{Dih}_{r})\circ_{S_{2}}\mathsf{e}(\mathsf{Cat})=\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\left(\eta_{k,d}(r)\frac{(1+2\dot{\mathscr{A}}_{0})^{2}}{(1-\psi_{2}(\mathscr{A}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathscr{A}_{0}^{\prime\prime}))^{k/j}}\right).$
@@
We now conclude by indicating how Theorems A and B follows from Theorem 6.4: in [23] (with a typo corrected in [54, Page 1]), Getzler proves
@@
7. Calculations
Calculations of $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d))$ for small $d$ and $n$, performed in SageMath [^5] The Sage notebook used to produce these calculations can be downloaded here. using Theorem A, are found in Tables 1, 2, and 3. The first step in producing these tables is to compute $\mathsf{b}_{1,r}^{\mathrm{nrt}}$ as follows. First, Theorem 4.1 gives the formula